OpenStudy (anonymous):

x^2+12x+36 perfect square ?

OpenStudy (anonymous):

i put (x+6)(x-6) but didnt work out

jimthompson5910 (jim_thompson5910):

x^2+12x+36 x^2+6x+6x+36 ... Note: 6 and 6 add to 12 and multiply to 36 (x^2+6x)+(6x+36) x(x+6)+6(x+6) (x+6)(x+6) (x+6)^2 So it is a perfect square

OpenStudy (anonymous):

oh, i thought i had to have one + and one minus? is that not the case?

jimthompson5910 (jim_thompson5910):

(x+6)(x-6) gives you x^2-36 (it's a difference of squares)

OpenStudy (anonymous):

(x+6)(x-6) = x^2 - 36 Difference of squares. Not the expression you have in this case.

jimthompson5910 (jim_thompson5910):

you're thinking of a difference of squares, which is binomial (ie has two terms)

OpenStudy (anonymous):

okay phew. got cha :)

OpenStudy (anonymous):

(a+b)^2 = a^2 + 2ab + b^2 (a-b)^2 = a^2 - 2ab + b^2 (a+b)(a-b) = a^2 - b^2

OpenStudy (anonymous):

thanks ^_^